OpenStudy (anonymous):

simplify.-7XY^2/28X^5_Y

OpenStudy (anonymous):

can some help me with this

OpenStudy (jackellyn):

-7xy^2/28x^5y When you divide variables, the exponents subtract. Thus, for the x exponent 1-5=-4 for the y exponent 2-1=1 for the coefficients -7/28=-1/4 -1/4(x^-4)y Following this form: a^(-b)= 1/(a^b) So (-1/4)y/x^4